4 years ago my father age was five times my age. at present the sum of our age is 50. what is my fathers present age? sir pls give me its step also?

Answer by TimothyLamb(4379)

---
x = father
y = son
---
x - 4 = 5(y - 4)
x + y = 50
---
put the system of linear equations into standard form:
---
x - 4 = 5y - 20
x + y = 50
---
x - 5y = -16
x + y = 50
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution by matrix method:
x = father = 39
y = son = 11
---
now by elimination:
---
( x - 5y = -16 ) * -1
x + y = 50
---
-x + 5y = 16
x + y = 50
---
add the system:
---
6y = 66
y = 11
---
x + 11 = 50
x = 39
---
solution by elimination method:
x = father = 39
y = son = 11
---
now by substitution:
---
x - 5y = -16
y = 50 - x
---
x - 5(50 - x) = -16
x - 250 + 5x = -16
6x = 234
x = 234/6
x = 39
---
y = 50 - 39
y = 11
---
solution by substitution method:
x = father = 39
y = son = 11
---
the methods agree !
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php